# DREAM: Dynamic Red-teaming across Environments for AI Models

### v1

@@ 4.1 Formalizing Attack Chain Generation
where the core of our DREAM engine lies in the concrete implementations of the state space $\mathcal{S}$, the action space $\mathcal{A}$, the belief update function $\tau$, and the heuristic value function $V$, which we detail in the subsequent sections.
To aid the reader in navigating our formal framework, we summarize the key symbols and notation used throughout our methodology in Table I.
TABLE I: Summary of key symbols and notation used in DREAM.
Symbol Description
PO-MDP Formulation (Sec 3.1)
$\langle\mathcal{S},\mathcal{A},\mathcal{O},P,R,\gamma\rangle$ The tuple defining the Partially Observable Markov Decision Process (PO-MDP).
$\mathcal{S}$ The latent state space of the target system.
$\mathcal{A}$ The action space, consisting of a discrete set of atom attacks.
$\mathcal{O}$ The observation space (natural language responses).
$P$ State transition probabilities (implicit).
$R$ The reward function.
$\gamma$ The discount factor.
$\pi$ A policy mapping belief states to actions.
$\pi^{*}$ The optimal policy.
$\pi_{\text{cond}}$ The heuristic policy of the Conductor agent (implemented by C-GPS).
$V(b_{t},a)$ The heuristic value function for action $a$ in belief state $b_{t}$.
State, Action, and Chain Dynamics
AA An attack chain, $A=(a_{1},\dots,a_{T})$.
$T$ The total number of timesteps (length) of an attack chain.
$t$ The current timestep.
$a_{t}$ The action (atom attack) selected at timestep $t$.
$b_{t}$ The Conductor’s belief state at timestep $t$ (implemented by the Unified World Model).
$o_{t+1}$ The observation received after executing action $a_{t}$.
$\tau$ The deterministic belief update function (Eq. 1).
$\mathcal{C}_{t}$ The dynamically generated candidate set of actions at timestep $t$.
Component Definitions (Sec 3.2)
$a$ An atom attack, a structured object in $\mathcal{A}$.
$E$ A target environment, $E\in\mathcal{E}$.
$\mathcal{E}$ The set of all distinct environments.
$\mathcal{K}_{t+1}$ Newly discovered entities extracted from observation $o_{t+1}$.
$\text{ctx}_{t+1}$ Context provisioned for the next action $a_{t+1}$.
$s(a_{t})$ A binary indicator of the success of action $a_{t}$.
$C_{\text{penalty}}$ A constant penalty applied for a failed action.
Value Function & Evaluation (Sec 3.3 - 3.4)
$R(b_{t},a_{t})$ The immediate reward for action $a_{t}$ (Eq. 4).
$Score(a)$ The pre-computed Atomic Score of an atom attack $a$.
$\text{req}(a)$ The set of prerequisite entities required for action $a$.
$\text{keys}(b_{t})$ The set of entities currently present in the belief state $b_{t}$.
$\text{EntityUsage}(b_{t},a)$ A metric quantifying how an action $a$ uses existing belief $b_{t}$ (Eq. 5).
$\text{Score}(A)$ The final score (discounted cumulative reward) for an entire attack chain $A$ (Eq. 6).
4.2 Defining PO-MDP: Action Space, States, and Rewards
4.2.1 Action Space $\mathcal{A}$: Multi-Agent Atom Attack Design
At the core of our framework is a rich and diverse action space $\mathcal{A}$. Each action $a\in\mathcal{A}$ is a structured object called an atom attack, which contains all the necessary information for execution. An atom attack is defined by key components: a description for semantic retrieval, and an action block that specifies tactical details, including a target environment $E\in\mathcal{E}$, a prompt template, and a set of entity requirements.
To construct this library, we design a systematic generation pipeline based on a multi-agent conceptual framework. This framework abstracts complex adversarial behaviors into three strategic roles, simulating a coordinated red team operation: ❶ Scout: A reconnaissance agent aiming to reduce uncertainty in the belief state $b_{t}$ by discovering new entities and information about the system’s latent state $\mathcal{S}$. ❷ Seeder: A state manipulation agent altering the latent state $\mathcal{S}$ to create vulnerabilities or preconditions for subsequent attacks. ❸ Exploiter: An agent exploiting the current belief state $b_{t}$ to achieve a malicious outcome, thus earning a high immediate reward.
These roles are designed to work together within a cohesive framework. The multi-agent abstraction serves as a blueprint for a powerful Large Language Model (LLM), which we instructed to generate a variety of atom attacks covering all three roles for each environment. This semi-automated process led to the creation of a comprehensive library of 1,986 atom attacks, spanning 349 distinct environments, forming a challenging and expansive action space for our Conductor agent.
4.2.2 State Space $\mathcal{S}$: Building A Unified Cross-Environment Adversarial Knowledge Graph
A major challenge in modeling long-chain attacks is overcoming the “information silos” typically found in multi-environment simulations. To address this, we introduce the Unified World Model, which serves as the concrete implementation of the Conductor’s belief state $b_{t}$. This model acts as a universal state reasoner, designed to integrate heterogeneous information from all environments into a coherent global state. At its core, this evolving model forms the foundation of our dynamic cross-environment adversarial knowledge graph (CE-AKG).
We realize this theoretical concept through the Unified Sandbox, a novel LLM-driven component that manages the state space $\mathcal{S}$ and approximates the state transition dynamics. The UnifiedSandbox maintains the belief state $b_{t}$ as a dynamic knowledge structure—specifically, a JSON object that represents a set of typed entities. Its core innovation is using the LLM to perform state inference, enabling the belief update function $\tau$ (as described in Eq. (1)) through two key operations:
Information Fusion When a natural language observation $o_{t+1}$ is received after an action, the UnifiedSandbox utilizes its LLM to parse the text, extract newly discovered entities $\mathcal{K}_{t+1}$, and intelligently merge them into the existing belief state $b_{t}$. This update results in the new belief state $b_{t+1}=b_{t}\cup\mathcal{K}_{t+1}$, ensuring that knowledge gained in any environment enriches the global world model.
Context Provisioning During the planning phase for the next action $a_{t+1}$, the UnifiedSandbox again leverages the LLM. Based on the action’s entity_requirements, it queries the global belief state $b_{t+1}$ to extract and format the required context $\text{ctx}_{t+1}$. This context is then used to instantiate the action’s prompt_template, equipping the next attack with all relevant information previously acquired.
This LLM-powered state management establishes an “information bridge” across environments. For example, a customer_id discovered by a Scout in a MissionControl environment is fused into the Unified World Model. This entity can later be provisioned to support a targeted Exploiter attack in a seemingly unrelated RetailFraud environment. This mechanism is crucial for enabling the automated generation of logically coherent, cross-domain domino attacks.
4.2.3 Reward Function $R$: Quantifying Immediate Utility
To guide the Conductor’s policy, we define a quantitative evaluation framework based on the reward function $R(b_{t},a_{t})$. A key component of this function is the Atomic Score, denoted as $Score(a)$, which represents the intrinsic potential of an atom attack. This score is pre-computed through large-scale, automated benchmarking of each attack’s success rate and impact.
Let $s(a_{t})$ be a binary indicator of the success of action $a_{t}$ at time step $t$, determined by an LLM. The immediate reward is then defined as:
$R(b_{t},a_{t})=s(a_{t})\cdot Score(a_{t})-(1-s(a_{t}))\cdot C_{\text{penalty}},$ (4)
where $C_{\text{penalty}}$ is a constant penalty applied when the action fails. This reward structure quantifies the immediate utility of each action and provides the basis for the strategic planning engine described in the next section.
4.3 The DREAM Engine: C-GPS for Strategic Planning
Having established the foundational components of our PO-MDP, we now introduce the DREAM engine, which implements the Conductor’s policy, $\pi_{\text{cond}}$, using a novel algorithm we call Contextualized Guided Policy Search (C-GPS). The primary challenge of C-GPS is navigating the expansive decision space to form long, impactful attack chains. It achieves this through two key capabilities: advanced state perception using the CE-AKG, and strategic long-term planning that drives a cascading domino effect.
4.3.1 State Perception via CE-AKG
In our framework, an observation $o_{t}$ is not just a local fact but a piece of global intelligence. The DREAM engine’s first task is to interpret this intelligence using the Unified World Model, enabling a deep level of situational awareness that spans across environments. This is achieved through the two main functions of the UnifiedSandbox, as described in Section 4.2.2.
First, Information Fusion converts a raw, unstructured natural language observation $o_{t}$ into a structured belief update $\mathcal{K}_{t}$, allowing the Conductor to continuously refine its understanding of the world state. Second, Context Provisioning enables the Conductor to leverage its global belief $b_{t}$. Before selecting an action, the engine queries the Unified World Model to provide contextually relevant information for potential attacks, no matter the source.
Powered by the CE-AKG, this mechanism transforms the Conductor’s perception. For example, an observed customer_id in the MissionControl environment becomes part of the global state through the Unified World Model. This global knowledge then informs the execution of a targeted attack in the RetailFraud environment. This ability to perceive states across environments is crucial for the engine’s success in identifying and exploiting systemic vulnerabilities.
4.3.2 C-GPS Policy: Heuristic Planning for Long Chains
To achieve a domino effect, the Conductor’s policy must prioritize actions that contribute to a long-term malicious objective over those that offer only myopic rewards. We steer the C-GPS algorithm using a heuristic value function, $V(b_{t},a)$, which holistically evaluates the strategic desirability of any given action $a$ from the current belief state $b_{t}$. The function is a weighted sum of three strategic components: ❶ Intrinsic Potential. This component is given by the action’s a priori Atomic Score, $Score(a)$, representing its inherent impact and likelihood of success. ❷ Information Exploitation. To explicitly encourage the formation of causal chains, we introduce an $\text{EntityUsage}(b_{t},a)$ term. This metric quantifies the degree to which an action leverages information already present in the belief state:
@@ 4.3.2 C-GPS Policy: Heuristic Planning for Long Chains
where $\text{req}(a)$ is the set of prerequisite entities for action $a$. A high EntityUsage score indicates that an action is a logical continuation of previous steps, effectively “locking” the dominoes together. ❸ Strategic Advancement. A strategic bonus is granted to actions that represent a significant tactical progression. This includes landmark achievements such as successful lateral movement to a new environment for the first time or achieving a privilege escalation, incentivizing the engine to explore and break new ground.
By balancing these three considerations, the value function $V(b_{t},a)$ provides a robust heuristic that guides the Conductor’s search for opportunistic, logically coherent, and high-impact attack sequences.
4.4 Attack Chain Generation & Evaluation
This section synthesizes the components defined above into an end-to-end workflow for generating and evaluating attack chains. We first detail the dynamic pipeline that the DREAM engine follows and then present the quantitative framework for assessing the overall efficacy of the generated attack sequences.
4.4.1 Dynamic Planning & Execution Pipeline
The Conductor agent operates in a feedback-driven loop, dynamically planning and adapting its strategy at each timestep $t$. This entire pipeline is driven by our C-GPS algorithm, which iteratively queries and updates the CE-AKG. As visualized in Fig. 2, this process consists of the following four stages:
Stage 1: Candidate Generation
The Conductor begins by narrowing the vast action space $\mathcal{A}$ to a small set of high-potential candidates, $\mathcal{C}_{t}$. This is achieved through a two-step process: first, a retrieval-augmented search over the Atom Attack Library uses the current belief state $b_{t}$ to find semantically relevant attacks. Second, these candidates are clustered to identify coherent strategic vectors, and the most promising cluster is selected to populate $\mathcal{C}_{t}$.
Stage 2: Action Selection
Next, the Conductor evaluates each candidate action $a\in\mathcal{C}_{t}$ using the strategic heuristic value function $V(b_{t},a)$ detailed in Section 4.3.2. It then selects the action $a_{t}$ that maximizes this value, in accordance with the policy $\pi_{\text{cond}}$ from Eq. (3).
Stage 3: Execution & Belief Update
The selected action $a_{t}$ is executed, yielding a natural language observation $o_{t+1}$. The UnifiedSandbox then parses this observation, updates the belief state to $b_{t+1}$ via Information Fusion, and computes the immediate reward $R(b_{t},a_{t})$.
Stage 4: Iteration & Backtracking
The process repeats from Stage 1 with the updated belief state. This loop structure allows the Conductor to continuously adapt its plan based on the outcomes of its actions. Crucially, the underlying decision process is a tree search. If a chosen path of actions leads to repeated failures, the engine has the capability to backtrack to a previous decision node and explore an alternative action from the candidate set, ensuring robustness in its pursuit of a successful attack chain.
4.4.2 Quantitative Evaluation of Attack Chains
To assess the effectiveness of a generated attack chain $A=(a_{1},\dots,a_{T})$, we use a final score to quantify its overall efficacy. As defined in Section 4.2.3, the immediate reward for each step, $R(b_{t-1},a_{t})$, reflects the action’s success and impact. The final score for the entire chain, $\text{Score}(A)$, is the discounted cumulative reward achieved by the policy:
$\text{Score}(A)=\sum_{t=1}^{T}\gamma^{t-1}R(b_{t-1},a_{t}),$ (6)
where $\gamma\in[0,1]$ is the discount factor, prioritizing efficiency by rewarding shorter paths to the objective. This scoring model encourages policies that generate successful, high-impact actions and penalizes failures. Ultimately, the final score serves as an objective, reproducible metric for benchmarking the robustness of different Target Agents against dynamic attacks generated by our framework.
4.5 The End-to-End Pipeline
This section synthesizes the previously described components into a cohesive, end-to-end workflow for generating and evaluating dynamic attack chains. The process begins with the offline, multi-agent generation of a comprehensive Atom Attack Library, which defines the action space for the Conductor agent. The Conductor then operates in a dynamic, feedback-driven loop. At each timestep, it leverages its current belief state (the CE-AKG) to generate a set of high-potential candidate actions. Guided by the Contextualized Guided Policy Search (C-GPS) policy, it selects the optimal action by maximizing a heuristic value function. This action is executed via the Unified Sandbox, which manages the stateful interaction with the target agent and returns an observation and reward. Based on this feedback, the Conductor updates its CE-AKG through information fusion if the action was successful, or triggers a backtracking mechanism to re-plan its strategy upon repeated failures. This iterative process allows the agent to construct long, causally-linked, and adaptive attack sequences. The complete pipeline is formally outlined in Algorithm 1.
@@ 4.5 The End-to-End Pipeline
Initialize: Belief state $b_{0}$ (CE-AKG), chain $A\leftarrow[\,]$
@@ 22     BacktrackToPreviousNode();
28Final Score: $\text{Score}(A)\leftarrow\sum_{t=1}^{|A|}\gamma^{t-1}R_{t}$;
@@
30Part 3: The Sandbox – Environment Interaction;
@@
Procedure: Sandbox.Execute($a,b$)
32
33prompt $\leftarrow$ ProvisionContext($a,b$);
@@
34
[…]
@@ 5.1 Experimental Setup
In this section, we detail the setup of our experiments, including the models we evaluate, the configuration of our automated attack framework, and the metrics used for evaluation.
5.1.1 Models Under Evaluation
Our evaluation encompasses a diverse suite of 12 state-of-the-art Large Language Models to ensure a comprehensive and generalizable assessment of agent safety. We selected models from two primary categories: leading proprietary, closed-source systems and prominent open-source alternatives. This selection allows for a robust comparison of safety postures across the LLM landscape.
The proprietary models evaluated are gemini-2.5-flash-nothinking, o4-mini, gemini-2.5-flash, GPT-5, gemini-2.5-pro, grok-4, and claude-sonnet-4.5-20250929-thinking. The open-source cohort consists of qwen3-235b-a22b, Kimi-K2-0711-preview, llama3.1-70b, qwen2.5-72b, and deepseek-v3.1-think-250821. To maintain consistency, all target agents were configured for deterministic output using greedy decoding.
@@ 5.1.2 Evaluation Metrics
To ensure a standardized and rigorous assessment, our metric design strictly aligns with the taxonomy proposed in SafetyBenchmark [yin2025safeagentbenchbenchmarksafetask]. We evaluate agent safety across 8 distinct risk categories and 9 specific failure modes, which serve as the fundamental granularity for our analysis.
Risk Categories Taxonomy. We adopt a comprehensive risk taxonomy covering the CIA triad (Confidentiality, Integrity, Availability) and broader safety concerns. As detailed in Table II, these categories range from data leakage to physical harm, ensuring a holistic safety assessment.
TABLE II: The 8 Risk Categories adopted from SafetyBenchmark used for classifying atom attacks.
@@
R3 Spread Misinfo Spread unsafe information/misinformation
@@
R6 Availability Compromise availability (DoS, resource exhaustion)
R7 Vulnerable Code Contribute to harmful/vulnerable code
R8 Produce Misinfo Produce unsafe information/misinformation
Failure Modes Taxonomy. Beyond outcome-based risks, we evaluate the cognitive root causes of safety failures. Table III lists the 9 failure modes, categorizing errors in reasoning, tool usage, and risk perception.
TABLE III: The 9 Failure Modes representing specific breakdowns in agent reasoning and tool usage.
@@
To move beyond simple success/failure rates, we further employ a five-dimensional evaluation framework that deconstructs an agent’s defense capabilities based on these taxonomies, as detailed in Table IV.
Overall Defense Score
This is the primary comprehensive metric for ranking agent safety, where a higher score indicates better defensive performance. It is derived from the final discounted cumulative reward of the attack chain, $\text{Score}(A)$ (Eq. (6)), reflecting the agent’s ability to successfully thwart high-impact actions and minimize overall harm throughout the entire adversarial campaign.
Attack Success Rate (%)
This metric quantifies the proportion of individual atom attack steps that successfully breached the agent’s defenses. It serves as a direct measure of the agent’s per-step resilience, with a lower rate being desirable.
Damage Mitigation
This metric evaluates an agent’s ability to limit the negative consequences of a successful breach. A high score indicates that even when compromised, the agent successfully contained the extent of the damage, for instance, by refusing the most harmful parts of a request or by sanitizing sensitive data before release.
Attack Observability
This metric measures the agent’s capacity to explicitly recognize and articulate potential threats. A high score signifies that the agent frequently identifies malicious requests and references its safety protocols in its responses, demonstrating a strong internal safety awareness, even if it is ultimately unable to prevent the exploit.
Contextual Isolation
This crucial metric assesses an agent’s resilience against the core threat of cross-environment attacks. It quantifies the agent’s ability to maintain strict boundaries between different operational contexts, preventing information gained in one environment from being leveraged to compromise another. A high score indicates robustness against “contextual fragility”, a key vulnerability our work identifies.
TABLE IV: Comprehensive Evaluation of LLM Agent Defenses Against Dynamic Adversarial Chains. ‘Overall Defense Score’ is the comprehensive safety score ($\uparrow$ higher is better). ‘Attack Success Rate’ represents the proportion of successful breaches ($\downarrow$ lower is better). The subsequent columns deconstruct the defense capabilities into specific dimensions. Models within each category are sorted by Overall Defense Score in ascending order.
Model Overall Defense Attack Success Damage Mitigation Attack Observability Contextual Isolation
Score $\uparrow$ Rate (%) $\downarrow$ $\uparrow$ $\uparrow$ $\uparrow$
@@
gemini-2.5-flash-nothinking 37.34 82 56.32 44.80 41.37
o4-mini 39.04 75 57.05 47.49 46.53
gemini-2.5-flash 40.89 76 60.27 47.41 44.90
GPT-5 45.39 62 62.85 52.39 52.00
gemini-2.5-pro 45.99 72 64.99 51.49 49.75
grok-4 61.17 43 75.32 64.68 67.53
claude-sonnet-4.5-20250929-thinking 67.30 24 76.80 71.95 84.23
@@
qwen3-235b-a22b 38.73 83 58.10 44.97 37.82
Kimi-K2-0711-preview 39.50 77 59.61 47.07 46.40
llama3.1-70b 41.40 76 58.43 48.21 43.91
qwen2.5-72b 43.37 73 61.61 48.65 44.99
deepseek-v3.1-think-250821 48.10 67 65.49 54.39 49.48

### v2

@@ 3.3.2 Heuristic Value Function
where $\mathcal{C}_{t}\subseteq\mathcal{A}$ is a dynamically generated candidate set (described in Section 3.3.3), and $V(b_{t},a)$ is a heuristic value function balancing three strategic considerations:
$\begin{split}V(b_{t},a)&=w_{1}\cdot\text{Score}(a)+\\ &\qquad w_{2}\cdot\text{EntityUsage}(b_{t},a)+w_{3}\cdot\text{StrategicBonus}(a),\end{split}$ (10)
(1) Intrinsic Potential: $\text{Score}(a)$ represents the action’s a priori impact and success likelihood (from the reward function in Section 3.2.3). (2) Information Exploitation: To explicitly encourage causal chaining, we define:
@@
where $\text{req}(a)$ is the set of required entities for action $a$. High EntityUsage indicates the action is a logical continuation of previous steps, effectively “locking dominoes together” in the attack chain.
(3) Strategic Advancement: A bonus term rewards landmark achievements such as first-time lateral movement to new environments or privilege escalation, incentivizing exploration and breakthrough moments.
This multi-faceted value function enables the Conductor to identify actions that are simultaneously high-impact, contextually coherent, and strategically progressive, which are the hallmarks of sophisticated multi-step exploits.
3.3.3 Dynamic Candidate Generation
To manage the large action space efficiently, we narrow $\mathcal{A}$ to a small candidate set $\mathcal{C}_{t}$ (typically $|\mathcal{C}_{t}|\approx 20{-}30$) through a two-stage process:
Stage 1: Retrieval-Augmented Search. Using the current belief state $b_{t}$ as a query, we perform semantic search over the Atom Attack Library to retrieve attacks relevant to current system knowledge and attack progress.
Stage 2: Strategic Clustering. Retrieved attacks are clustered by strategic intent (e.g., “database exfiltration”, “lateral movement to cloud infrastructure”). The Conductor selects the most promising cluster based on attack progress and unexplored vectors, populating $\mathcal{C}_{t}$ with diverse yet strategically coherent options.
This candidate generation mechanism balances the exploration of diverse attack vectors with the exploitation of promising strategic directions, enabling efficient navigation of the vast action space.
3.4 End-to-End Attack Chain Generation
@@
Init: Belief state $b_{0}$ (CE-AKG), chain $A\leftarrow[\,]$
@@ 22     Backtrack to Previous Node();
28 end for
@@
30Final Score: $\text{Score}(A)\leftarrow\sum_{t=1}^{|A|}\gamma^{t-1}R_{t}$;
@@
32Part 3: The Sandbox – Environment Interaction;
33
34Procedure Sandbox.Execute($a,b$)
35
36 prompt $\leftarrow$ ProvisionContext($a,b$);
@@
37
38 $o\leftarrow\text{TargetAgent.query}(\text{prompt})$;
@@ 40 $(R,s)\leftarrow\text{EvaluateOutcome}(o,a)$;
We now synthesize all components into a complete workflow. Algorithm 1 presents the formal procedure; here we describe the four-stage iterative loop:
Stage 1: Candidate Generation. The Conductor queries the Atom Attack Library using its current belief state $b_{t}$, retrieving and clustering semantically relevant attacks to form the candidate set $\mathcal{C}_{t}$ (Section 3.3.3).
Stage 2: Action Selection. Each candidate $a\in\mathcal{C}_{t}$ is evaluated using the heuristic value function $V(b_{t},a)$. The Conductor selects $a_{t}=\pi_{\text{cond}}(b_{t})$ that maximizes this value (Equation 9).
Stage 3: Execution & Belief Update. The Unified Sandbox executes $a_{t}$ against the target agent, receiving observation $o_{t+1}$. It then: (1) performs information fusion to update the CE-AKG ($b_{t+1}=\tau(b_{t},a_{t},o_{t+1})$), and (2) computes immediate reward $R(b_{t},a_{t})$ via success detection and atomic score lookup.
Stage 4: Iteration & Backtracking. The process repeats from Stage 1 with updated belief $b_{t+1}$. Crucially, the underlying decision process maintains a search tree: if repeated failures occur along a path, the engine backtracks to a previous decision node and explores alternative candidates, ensuring robustness.
This feedback-driven loop enables the Conductor to construct attack chains that are: (1) adaptive, responding to target agent defenses in real-time; (2) causally coherent, leveraging information across environments; and (3) strategically deep, balancing immediate exploitation with long-term attack development.
3.4.1 Quantitative Evaluation Metrics
For a generated attack chain $A=(a_{1},\dots,a_{T})$, we compute the final score as the discounted cumulative reward:
$\text{Score}(A)=\sum_{t=1}^{T}\gamma^{t-1}R(b_{t-1},a_{t}),$ (12)
where $\gamma$ prioritizes efficiency by rewarding shorter paths to objectives. This metric provides an objective, reproducible measure of attack chain efficacy, allowing for the systematic comparison of target agent robustness across heterogeneous defense settings.
@@ 4 Experiments
This section describes our experimental setup: the evaluated models, the attack framework configuration, and the evaluation metrics.
@@ 4.1 Experimental Setup
4.1.1 Evaluated Agents
To ensure a representative assessment of the current LLM landscape, we evaluate a diverse suite of 12 state-of-the-art models, spanning both frontier proprietary systems and prominent open-source architectures. The proprietary group includes Gemini-2.5-Flash, o4-mini, GPT-5, Grok-4, and Claude-Sonnet-4.5. These are benchmarked against an open-source counterparts, which comprise Qwen2.5-72B, Qwen3-235B, Kimi-K2, Llama-3.1-70B, and DeepSeek-V3.1. To ensure reproducibility and eliminate stochastic variance, all agents were configured for deterministic execution via greedy decoding.
@@ 4.1.2 Evaluation Metrics
To ensure a standardized and rigorous assessment, our metric design builds upon the taxonomy proposed in SafetyBenchmark [43]. We evaluate agent safety across 8 distinct risk categories and 9 specific failure modes, which serve as the fundamental granularity for our analysis.
Risk Categories Taxonomy. We adopt a comprehensive risk taxonomy covering the CIA triad (Confidentiality, Integrity, Availability) and broader safety concerns. As detailed in Table 2, these categories range from data leakage to physical harm, ensuring a holistic safety assessment.
Table 2: Risk categories for atom attack classification. Eight categories adapted from SafetyBenchmark [43].
@@
R3 Spread Misinfo Spread unsafe info/misinformation
@@
R6 Availability Compromise availability (DoS)
R7 Vulnerable Code Contribute to harmful code
R8 Produce Misinfo Produce unsafe info/misinformation
Failure Modes Taxonomy. To extend our analysis beyond outcome-based risks, we evaluate the cognitive root causes of safety failures. Table 3 presents a taxonomy of 9 failure modes, categorizing errors across reasoning, tool usage, and risk perception.
Table 3: Agent failure modes. Nine categories of breakdowns in reasoning and tool usage.
@@
Comprehensive Defense Metrics. To move beyond simple binary success rates, we employ a five-dimensional framework to deconstruct defense capabilities:
  • Overall Defense Score ($\uparrow$): The primary safety ranking metric derived from the discounted cumulative reward $\text{Score}(A)$ (Eq. 12). It reflects the agent’s ability to thwart high-impact actions over the entire campaign.
  • Attack Success Rate (ASR, $\downarrow$): The proportion of individual atom attack steps that successfully breach the agent’s defenses. A lower rate indicates better per-step resilience.
  • Damage Mitigation ($\uparrow$): Evaluates the agent’s ability to limit negative consequences upon breach (e.g., by sanitizing data or refusing partial requests), ensuring “defense in depth.”
  • Attack Observability ($\uparrow$): Measures the agent’s capacity to explicitly recognize threats. A high score signifies that the agent identifies malicious intent and cites safety protocols, even if the attack partially succeeds.
  • Contextual Isolation ($\uparrow$): Assesses resilience against cross-environment threats. It quantifies the ability to maintain strict boundaries between contexts, preventing information gained in Environment A from compromising Environment B.
Table 4: Defense performance against dynamic adversarial chains. Overall Defense Score (composite safety metric, $\uparrow$) and Attack Success Rate ($\downarrow$) are reported. Models ranked by Overall Defense Score. Bold indicates best performance.The widespread high Attack Success Rates (>60% for most models) demonstrate DREAM’s effectiveness in uncovering systemic “contextual fragility” and long-chain vulnerabilities that remain hidden in traditional static benchmarks.
Model Overall Defense Attack Success Damage Attack Contextual
Score ($\uparrow$) Rate ($\downarrow$) Mitigation ($\uparrow$) Observability ($\uparrow$) Isolation ($\uparrow$)
@@
Gemini-2.5-Flash-NT 37.34 82.00 56.32 44.80 41.37
o4-mini 39.04 75.00 57.05 47.49 46.53
Gemini-2.5-Flash 40.89 76.00 60.27 47.41 44.90
GPT-5 45.39 62.00 62.85 52.39 52.00
Gemini-2.5-Pro 45.99 72.00 64.99 51.49 49.75
Grok-4 61.17 43.00 75.32 64.68 67.53
Claude-Sonnet-4.5 67.30 24.00 76.80 71.95 84.23
@@
Qwen3-235B-A22B 38.73 83.00 58.10 44.97 37.82
Kimi-K2-Preview 39.50 77.00 59.61 47.07 46.40
Llama-3.1-70B 41.40 76.00 58.43 48.21 43.91
Qwen-2.5-72B 43.37 73.00 61.61 48.65 44.99
DeepSeek-V3.1 48.10 67.00 65.49 54.39 49.48